OpenStudy (anonymous):

write in scientific notation 28.7731 answer 2.87731 x 10^-1?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

10^1 , NOT 10^-1

OpenStudy (ash2326):

it'd be 2.87731*10^1

OpenStudy (king):

yeah 10^1

OpenStudy (turingtest):

2.87731 x 10^-1=2.87731/10=0.287731 so that is wrong 2.87731 x 10^1=2.87731(10)=28.7731 which is correct